# Linearized Stability of Harada Thin-Shell Wormholes

### v1

@@ I Introduction
Traversable Lorentzian wormholes [1, 2] are known as solutions for Einstein’s field equations. The Conventional way to find a wormhole solution is to select some equations of state of matter such as Chaplygin gas [3], phantom energy [4, 5], and/or quintessence [6]. Then one can apply geometrical constructions such as thin-shell spacetimes [7], evolving wormholes [8], rotating spacetimes [9], dust shell wormholes [10], and/or Casimir wormholes [11] in order to alleviate the violation of energy conditions associated with the flaring-out condition that distinguishes wormhole solutions from other solutions. A thin-shell is a singular boundary hypersurface with energy-momentum conditions as stabilizers for such surface. If a thin shell is used to connect two separate spacetime with flaring-out condition, this thin shell forms a thin-shell wormhole. There are numerous studies that consider different black holes creating thin-shell wormholes in de-Sitter (dS) and anti-de-Sitter (AdS) spacetimes [12, 13, 14, 15, 16, 17]. Stability of these thin-shell wormholes are examined too [18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37].
@@
Recently, Harada targeted three main obstacles in Einstein’s general theory of relativity [38]. These obstacles are: i. the nature of the cosmological constant as an integration constant, ii. the derivation of conservation of energy-momentum tensor rather than being as conjectured an ad hoc, and iii. the usage of conformal flat metric as a vacuum solution and the unphysical solutions associated with it. Harada new gravity theory shows that accelerating expansion of the universe naturally appears as a consequence of the beyond general relativity gravitational field equations even in the absence of the conventional cosmological constant and/or dark energy [39]. Harada endeavors are considered a conformal Killing gravity theory in which the energy-momentum tensor is corrected by a divergenceless conformal Killing tensor, and thus the theory is compared to Cotton gravity [40]. Harada theory comes with many physical consequences such as generalized solutions of Schwarschild-Reissner-Nordström-AdS and regular black holes [41, 42], non-asymptotically flat traversable wormholes [43], black bounces [44], and pp-waves [45].
@@
In this letter we study thin-shell wormholes in Harada gravity theory. In section ($\mathbf{II}$), we use Visser’s cut-and-paste technique [47, 46], together with Darmois-Israel-(Sen) junction conditions [48, 49, 50, 51], to connect two conformal Killing gravity regions of spacetime through a thin shell. The cut-and-paste method with the junction conditions are comprehensively reviewed in Ref. [52]. The methodology of cut-and-paste provides the advantages of the employment of lesser amount of exotic matter, and hence traveling objects through this wormhole avoid the regions occupied by exotic matter [47]. This results in confining the exotic matter at the thin-shell regions similar to the matter studied in other examples of thin-shell wormholes [20, 53, 36, 54, 55, 56, 57, 58, 59, 60, 61, 62]. Also, we study the components of the hypersurface energy-momentum tensor using the extrinsic curvature components. We utilize these components to obtain the stress and pressure then comment on the violation of energy conditions because of the exotic matter at the wormhole throat. From the discontinuity in the extrinsic curvature, we compare the effect of extra fourth order term in Harada metric on the spacetime to the effect of the usual second order cosmological constant term on spacetimes studied in Ref. [63]. Additionally we comment on the physics of attraction and repulsion on the wormhole throat in terms of the acceleration.
@@ II Dynamics of Thin-Shell Wormhole
For Harada black hole with no radial pressure, $p_{r}=0$, and a mass density localized at the throat $\rho=\sigma_{0}\leavevmode\nobreak\ \delta(r-R_{0})$, the total amount of exotic matter necessary to keep the wormhole open is
@@ III Linearized Stability Analysis
Next, we rearrange eq.(14) and define a potential function
@@
Figure 4: Regions of stability of the thin-shell wormhole for the Bardeen de-Sitter solution for fixed values of $M=1$ and $\Lambda=10^{-2}$, and different values of $N$. Stable regions are the blue shaded domains.
@@ IV Discussion
After that, we analyze the linearized stability of Harada thin-shell wormhole by examining the concavity behavior on the “speed of sound” as a function in Harada parameter, the mass, and the cosmological constant. And we notice the change in stability regions upon varying Harada parameter while both mass and cosmological constant are constant. The analysis is demonstrated in fig.(4). We conclude that for a lesser value of cosmological constant and lesser/larger value of magnetic charge, relative to the amount of mass, regions of stability vary. Once the mass is equal to the Harada parameter, we no longer have stability regions. Therefore in the Harada spacetime, and for a small value of the cosmological constant, thin-shell wormholes are kept open as long as the value of Harada parameter is away from the value of mass.

### v2

@@ I Introduction
Traversable Lorentzian wormholes [1, 2] are known as solutions for Einstein’s field equations. The Conventional way to find a wormhole solution is to select some equations of state of matter such as Chaplygin gas [3], phantom energy [4, 5], and/or quintessence [6]. Then one can apply geometrical constructions such as thin-shell spacetimes [7], evolving wormholes [8], rotating spacetimes [9], dust shell wormholes [10], and/or Casimir wormholes [11] in order to alleviate the violation of energy conditions associated with the flaring-out condition that distinguishes wormhole solutions from other solutions. A thin-shell is a singular boundary hypersurface with energy-momentum conditions as stabilizers for such surface. If a thin shell is used to connect two separate spacetime with flaring-out condition, this thin shell forms a thin-shell wormhole. There are numerous studies that consider different black holes creating thin-shell wormholes in de-Sitter (dS) and anti-de-Sitter (AdS) spacetimes [12, 13, 14, 15, 16, 17]. Stability of these thin-shell wormholes are examined too [18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37]. Also, one should consider several other examples of wormhole solutions that do not require exotic matter such as those supported by the Dirac-Maxwell fields [38, 39] and those supported by applying different geometrical modifications to the usual spacetime [40, 41, 42, 43, 44, 45]. It is worth noting that the stability of wormholes is an old issue in theories of gravity, and until this moment there is no wormhole solution for which stability has been realized under general perturbations, see the examples in Ref. [46, 47, 48] and the cited references there.
@@
Recently, Harada targeted three main obstacles in Einstein’s general theory of relativity [49]. These obstacles are: i. the nature of the cosmological constant as an integration constant, ii. the derivation of conservation of energy-momentum tensor rather than being as conjectured an ad hoc, and iii. the usage of conformal flat metric as a vacuum solution and the unphysical solutions associated with it. Harada new gravity theory shows that accelerating expansion of the universe naturally appears as a consequence of the beyond general relativity gravitational field equations even in the absence of the conventional cosmological constant and/or dark energy [50]. Harada endeavors are considered a conformal Killing gravity theory in which the energy-momentum tensor is corrected by a divergenceless conformal Killing tensor, and thus the theory is compared to Cotton gravity [51]. Harada theory comes with many physical consequences [52, 53] such as generalized solutions of Schwarschild-Reissner-Nordström-AdS and regular black holes [54, 55], non-asymptotically flat traversable wormholes [56], black bounces [57], and pp-waves [58].
@@
We are interested in studying the stability of thin-shell wormholes in Harada gravity theory for the following two main reasons. The first one is that finding a stable soltuion to traversable wormholes in conformal Killing gravity is a prominent result because such solution is not allowed by the null energy condition in the context of the conventional general theory of relativity [59]. The second reason is that there is a class of wormhole solutions of conformal Killing gravity found in Ref. [56] the does not demand the support of exotic matter or dark energy to stabilize the wormholes; and this stimulates considering the thin-shell wormholes within the same theory. In section ($\mathbf{II}$), we use Visser’s cut-and-paste technique [61, 60], together with Darmois-Israel-(Sen) junction conditions [62, 63, 64, 65], to connect two conformal Killing gravity regions of spacetime through a thin shell. The cut-and-paste method with the junction conditions are comprehensively reviewed in Ref. [66]. The methodology of cut-and-paste provides the advantages of the employment of lesser amount of exotic matter, and hence traveling objects through this wormhole avoid the regions occupied by exotic matter [61]. This results in confining the exotic matter at the thin-shell regions similar to the matter studied in other examples of thin-shell wormholes [20, 67, 36, 68, 69, 70, 71, 72, 73, 74, 75, 76]. Also, we study the components of the hypersurface energy-momentum tensor using the extrinsic curvature components. We utilize these components to obtain the stress and pressure then comment on the violation of energy conditions because of the exotic matter at the wormhole throat. From the discontinuity in the extrinsic curvature, we compare the effect of extra fourth order term in Harada metric on the spacetime to the effect of the usual second order cosmological constant term on spacetimes studied in Ref. [77]. Additionally we comment on the physics of attraction and repulsion on the wormhole throat in terms of the acceleration.
@@ II Dynamics of Thin-Shell Wormhole
For Harada black hole with no radial pressure and a mass density localized at the throat $\rho=\sigma_{0}\leavevmode\nobreak\ \delta(r-R_{0})$, the total amount of exotic matter necessary to keep the wormhole open is
@@ III Linearized Stability Analysis
Next, rearrange eq.(14) and define a potential function
@@
Figure 4: Regions of stability of the thin-shell wormhole for the Harada solution with fixed values of $M=1$ and $\Lambda=10^{-2}$, and different values of $N$. Stable regions are the blue shaded domains.
@@ IV Discussion
After that, we analyze the linearized stability of Harada thin-shell wormhole by examining the concavity behavior of eq.(28), that of the “speed of sound”, as a function in Harada parameter, the mass, and the cosmological constant. And we notice changes in stability regions upon varying Harada parameter while keeping the mass and cosmological constant fixed. Besides being considered as a stability measure for the wormhole solutions, the speed of sound function in principle can help detecting such systems by revealing the thermodynamics of the equations of states of the ordinary and/or exotic matter supporting the wormhole system. The analysis of stability regions of speed of sound is demonstrated in fig.(4), and we conclude that for a lesser value of cosmological constant and lesser/larger value of magnetic charge, relative to the amount of mass, regions of stability vary. Once the mass is equal to the Harada parameter, we no longer have stability regions. Therefore in the Harada spacetime, and for a small value of the cosmological constant, thin-shell wormholes are kept open as long as the value of Harada parameter is away from the value of mass.